OpenStudy (anonymous):

more algebra 2 word problems heathers age is 3 years more than one half her cousin rachels age. if the difference in their ages is less than 4 years, what is the oldest that rachel can be in years?

OpenStudy (anonymous):

Let the Heather's age be : x And the Rachel's Age be : y. So: \[x = 3 + \frac{1}{2}y\] \[2x = 3 + y\] \[2x - y = 3\] \[x - y < 4\] \[x - 2x + 3 < 4\] \[-x < 1 \implies x > -1\]

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

\[3 + \frac{1}{2}y - y < 4 \implies 3 - \frac{1}{2}y<4 \implies 6-y< 8 \implies y> -2\]

OpenStudy (anonymous):

Why thank you???

OpenStudy (anonymous):

you helped a lot

OpenStudy (anonymous):

I am solving it wrong.. Ha ha ha.. Age cannot be negative..

OpenStudy (anonymous):

the question might be incomplete, but I will begin with \[r-h<4\] or \[h-r<4\] where \[h=3+\frac{ 1 }{ 2 }r\] this simply means that Rachel (r) is older than Heather (h) \[r>h\] by \[r=2h-3\] I will then pick the first inequality, r - h < 4 \[r-(3+\frac{ 1 }{ 2 }r)<4\]

OpenStudy (anonymous):

How r is older ???

OpenStudy (anonymous):

eathers age is 3 years more than one half her cousin rachels age

OpenStudy (anonymous):

So heather is older no ??

OpenStudy (anonymous):

h=\[\frac{ 1 }{ 2 }r+3\]

OpenStudy (anonymous):

Put r = 2 there...

OpenStudy (anonymous):

not twice, one-half

OpenStudy (anonymous):

For r = 2, h = 4 implies h is older..

OpenStudy (anonymous):

I was thinking that it should be h=2r+3, but that is incorrect if you put it in words.

OpenStudy (anonymous):

Yes this is incorrect..

OpenStudy (anonymous):

let me drink coffee and then give me 5 mins. I just woke up...

OpenStudy (anonymous):

\[h = 3 + \frac{1}{2}r\] This is according to question and h is older than r..

OpenStudy (dumbcow):

haha , @waterineyes , in your equations you were assuming heather is older but rachael is always older \[x = 3 + \frac{1}{2}y\] \[y -x < 4\] by substitution \[y-(3+\frac{1}{2}y) < 4\] \[\frac{1}{2}y < 7\] \[y < 14\] Thus oldest rachael can be is 14

OpenStudy (anonymous):

Their difference is less than 4 implies it can be 3, 2, 1 or it can be zero also.. x > -1, x can be 1, 2, 3, 4 etc y > -2, y can be 1, 2, 3, 4 etc But there is one condition also that x = 3 + 1/2y..

OpenStudy (anonymous):

Put y = 2 there..

OpenStudy (dumbcow):

sorry ...not always but we are looking at maximizing y

OpenStudy (anonymous):

Yeah this can be possibility @dumbcow Thanks for correcting me..

OpenStudy (anonymous):

There is something wrong with the question too.. The difference will be less than 4 and x = 3 + y/2 For y = 2 we get x = 4 Also x - y < 4 Both the condition satisfy..

OpenStudy (anonymous):

let us break the words down: Heather's age is 3 years more (+3) than one-half (1/2) of Rachel's age. Then think if this makes sense: Rachel is 3 years younger (-3) twice Heather's age 2h.

OpenStudy (anonymous):

r=2h-3

OpenStudy (anonymous):

\[x = 3 + \frac{1}{2}y \implies y = 2x - 3\] Rachel is 3 years less than twice the Heather's age..

OpenStudy (anonymous):

this is why I picked r and h, instead of x and y for simplicity

OpenStudy (anonymous):

My issue was not that.. I am just saying that Rachel will be younger than heather.. And yes you can take h and r for simplicity..

OpenStudy (anonymous):

How can a person at the age of 2 younger than heather by 2 years be older ?? @dumbcow

OpenStudy (anonymous):

Heather's age is 3 years more (+3) than one-half (1/2) of Rachel's age. Heather's age is best understood by looking that it is first half of Rachel's age. And then add 3 years.

OpenStudy (anonymous):

Yes but for r = 2 all situations fail..

OpenStudy (anonymous):

Are you really typing @panlac01

OpenStudy (anonymous):

\[r-(3+\frac{ 1 }{2 }r)<4\] \[r-\frac{ 1 }{ 2 }r<7\] \[\frac{ 1 }{ 2 }r <7\] \[r<14\]

OpenStudy (anonymous):

So, maximum \(r\) will be 13 years..

OpenStudy (anonymous):

No it will be 12 years..

OpenStudy (anonymous):

13 years and 11 months and 30 days

OpenStudy (anonymous):

Ha ha ha... For 13 years, you will get x as not an natural number.. I think r will be 12 years..

OpenStudy (anonymous):

but that's how math works. it doesn't care when a person is half. we assume that it is 1 because there can't be a half a person.

OpenStudy (anonymous):

It is basically when we assume x and y we assume that these are ages so ages can have only natural values.. For r = 13 you are getting h not a natural number so r = 13 is not possible..

OpenStudy (anonymous):

it is true that it is not a natural number, but we can say that Heather is 9 years old and 6 months at minimum

OpenStudy (anonymous):

and to go back to the question, Rachel is less than 14 years of age. Meaning, she is 13 years old, 11 months and 30 days old at most. If you round this of, you will get 14. So I think it is safe to say that Rachel is 13-years-old and you can never be wrong for making that as a final answer.